07.5x-0.9=5.7x+4.5
Answer by rapaljer(4551)

If you don't like those ugly decimal numbers, then multiply both sides of the equation by 10, which will clear the decimals. You will have:
75x - 9 = 57x + 45

Subtract 57x from each side:
75x - 57x - 9 = 57x - 57x + 45
18x - 9 = 45

Next, add + 9 to each side:
18x - 9 + 9 = 45 + 9
18x = 54

Divide both sides by 18:
x = 3

R^2 at SCC